# Can LLMs Handle WebShell Detection? Overcoming Detection Challenges with Behavioral Function-Aware Framework

### v4

@@ 6 Conclusion
This paper systematically demonstrates that baseline LLMs are suboptimal for WebShell detection, exhibiting a precision-recall trade-off that underperforms state-of-the-art methods. We propose the Behavioral Function-Aware Detection (BFAD) framework, which overcomes these limitations by focusing analysis on critical code segments and optimizing in-context learning. BFAD enables our LLM-based approach to surpass traditional benchmarks.
Appendix A Limitations and Future Work
Our research also highlights several limitations that present important directions for future investigation in this emerging field.
A.1 Dataset Generalization
A fundamental challenge in WebShell research is the scarcity of diverse, real-world data. Our reliance on public datasets, like that of prior work, limits the evaluation of model generalization against novel, sophisticated threats. To address this, future work should focus on using LLMs to generate large-scale synthetic benchmarks. A large-scale, LLM-generated dataset would provide a privacy-preserving and diverse alternative for stress-testing model robustness against a wider range of attack vectors.
A.2 Model and System Robustness
Our framework’s static analysis approach remains vulnerable to advanced obfuscation and zero-day attacks. Future work should therefore focus on enhancing model robustness and system adaptability. To counter evasion, models could integrate dynamic analysis by examining runtime function call traces from sandboxed execution. To deepen semantic understanding, domain-specific fine-tuning of code-centric LLMs on curated WebShell datasets is a crucial next step. For practical deployment, these enhanced models could be integrated into a multi-stage ”fast and slow” system, using a smaller model for rapid filtering and a larger one for in-depth analysis of high-risk samples. To ensure long-term resilience, this pipeline could be governed by an autonomous agent that continuously identifies new attack patterns and automatically updates the detection models.
Appendix B Prompt Details
@@
Appendix C Critical Function Details
@@
Appendix D Dataset Details

### v5

@@ 7 Conclusion
This paper provides a systematic study of LLMs for WebShell detection and shows that off-the-shelf prompting is not sufficient: across model scales, vanilla LLMs exhibit a pronounced precision–recall imbalance and fall short of strong supervised baselines. To bridge this gap, we introduce Behavioral Function-Aware Detection (BFAD), which adapts LLM inference to WebShell behaviors through (i) critical-function anchoring, (ii) context-aware extraction that fits long scripts into a limited prompt budget, and (iii) behavior-weighted demonstration retrieval via WBFP. Empirically, BFAD consistently improves all evaluated LLMs and enables several models to match or surpass state-of-the-art detectors. Overall, our findings clarify both the failure modes and the promise of LLM-based WebShell detection, and suggest that behavior-centric prompting and retrieval are key to making LLMs reliable in security-critical code analysis.
Appendix A Prompt Details
@@
Appendix B Critical Function Details
@@
Appendix C Dataset Details